Solve for x: 5{2x-3}-{x+3}=0
Answer by kelsiesmurf(9)

5{2x-3}-{x+3}=0
10x - 15 - x - 3 = 0
9x - 18 = 0
9x = 18
x = 2
you can check this by plugging in 2
5{2*2 - 3} - {2 + 3} = 0
5{4 - 3} - 5 = 0
5 - 5 = 0